# MAPSS: Manifold-based Assessment of Perceptual Source Separation

### v1

@@ B.1 The PS Measure
$\left.\frac{\hat{B}_{i}^{(d)}\sqrt{\boldsymbol{r}_{i,i}^{T}\boldsymbol{S}_{i}^{-1}\boldsymbol{r}_{i,i}}+\hat{A}_{i}^{(d)}\sqrt{\boldsymbol{r}_{i,j^{*}}^{T}\boldsymbol{S}_{j^{*}}^{-1}\boldsymbol{r}_{i,j^{*}}}}{\left(\hat{A}_{i}^{(d)}+\hat{B}_{i}^{(d)}\right)^{2}}+\right.\left.L_{i}^{\textrm{PS}}\sqrt{\varepsilon^{\mathrm{PS}}\left(\widehat{A}_{i}^{(d)}\right)+\varepsilon^{\mathrm{PS}}\left(\widehat{B}_{i}^{(d)}\right)}\right\}\geq 1-\delta_{i}^{\textrm{PS}}.$
We now analyze the obtained expression separately for the deterministic and probabilistic terms. In the former term, the two square-root terms are energies that leak into the truncated complement after regressing out the retained $d$ diffusion coordinates. Intuitively, $\boldsymbol{\Sigma}^{(d)}_{j}$ encodes the local anisotropy of cluster $j$ in the kept coordinates, $\boldsymbol{C}_{j}$ represents coupling of residual energy in the truncated block, and $\boldsymbol{\Sigma}^{(\perp)}_{j}$ is the spread that remains in the truncated block. Therefore, larger $\boldsymbol{S}_{j}$ down-weights complement deviations, reducing the bias, while $\boldsymbol{r}_{i,j}$ and $\boldsymbol{S}_{j}$ co-vary through the cross-covariance $\boldsymbol{C}_{j}$. Namely, increasing $\boldsymbol{C}_{j}$ shrinks both $\boldsymbol{r}_{i,j}$ and $\boldsymbol{S}_{j}$, while decreasing $\boldsymbol{C}_{j}$ does the opposite. The practical rule is to prevent tiny $\lambda_{\min}\left(\boldsymbol{S}_{j}\right)$, e.g., by promoting such directions into the kept set via a local choice of $d$, and shape distortions so the complement is predictable from the kept coordinates, keeping $\boldsymbol{r}_{i,j}$ small.
For the probabilistic part, its width reflects uncertainty in the empirical centroid and covariance of the attributed and nearest foreign clusters, where $\lambda_{\max}\!\bigl{(}\widehat{\boldsymbol{\Sigma}}^{(d)}_{j}\bigr{)}$ and $n_{j,\mathrm{eff}}$ determine the width primarily. Practically, correlated distortions shrink $n_{j,\mathrm{eff}}$ and widen the bound, and a smaller spectral flatness ratio $r_{j}$ yields tighter matrix concentration. As expected, the probabilistic piece dominates the deterministic as shown in Figure 9, which underlines the importance of cluster construction and dependence control.
B.2 The PM Measure
As in the PS case, we start with the truncation error and assume access to cluster statistics. Let us reconsider Eqs. (70)–(82), but with two adjustments. First, the cluster coordinates are centered around the cluster reference embedding and not the cluster mean. Given m=Nd1m=N-d-1, we define:
$\boldsymbol{\Delta}^{(d)}_{i,p}$ $=\boldsymbol{\Psi}^{(d)}_{t}(\mathbf{x}_{i,p})-\boldsymbol{\Psi}^{(d)}_{t}(\mathbf{x}_{i,0})\in\mathbb{R}^{d},$ (107)
$\boldsymbol{\Delta}^{(\perp)}_{i,p}$ $=\boldsymbol{\Psi}^{(\perp)}_{t}(\mathbf{x}_{i,p})-\boldsymbol{\Psi}^{(\perp)}_{t}(\mathbf{x}_{i,0})\in\mathbb{R}^{m}.$ (108)
$\tilde{c}_{i}=c_{i}\setminus\Psi_{t}(\mathbf{x}_{i}).$ (109)
The cluster $\tilde{c}_{i}$ has difference $\boldsymbol{\Delta}_{i,p}\in\mathbb{R}^{N-1}$} for every $p\in\{1,\ldots,N_{p}\}$} and covariance $\boldsymbol{\tilde{\Sigma}}_{i}\in\mathbb{R}^{(N-1)\times(N-1)}$} that hold (28):
$\boldsymbol{\Delta}_{i,p}=\begin{bmatrix}\boldsymbol{\Delta}_{i,p}^{(d)}\\ \boldsymbol{\Delta}_{i,p}^{(\perp)}\end{bmatrix},\quad\boldsymbol{\tilde{\Sigma}}_{i}=\begin{bmatrix}\boldsymbol{\tilde{\Sigma}}_{i}^{(d)}&\boldsymbol{\tilde{C}}_{i}\\ \boldsymbol{\tilde{C}}_{i}^{T}&\boldsymbol{\tilde{\Sigma}}_{i}^{(\perp)}\end{bmatrix}$, (110)
with $\boldsymbol{\tilde{\Sigma}}_{i}^{(\perp)}\in\mathbb{R}^{m\times m}$ and $\boldsymbol{\tilde{C}}_{i}\in\mathbb{R}^{d\times m}$ being:
$\boldsymbol{\tilde{\Sigma}}_{i}^{(\perp)}=\frac{1}{\Big{|}\tilde{c}^{(\perp)}_{i}\Big{|}-1}\sum_{\boldsymbol{\psi}\in\tilde{c}_{i}^{(\perp)}}\left(\boldsymbol{\psi}-\Psi_{t}^{(\perp)}(\mathbf{x}_{i})\right)\left(\boldsymbol{\psi}-\Psi_{t}^{(\perp)}(\mathbf{x}_{i})\right)^{T},$ (111)
$\boldsymbol{\tilde{C}}_{i}=\frac{1}{\Big{|}\tilde{c}^{(\perp)}_{i}\Big{|}-1}\sum_{p=1}^{N_{p}}\left(\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i,p})-\Psi_{t}^{(d)}(\mathbf{x}_{i})\right)\left(\boldsymbol{\Psi}_{t}^{(\perp)}(\mathbf{x}_{i,p})-\Psi_{t}^{(\perp)}(\mathbf{x}_{i})\right)^{T}.$ (112)
In N1N-1 dimensions, the squared Mahalanobis distance from $\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,p})$ to $\tilde{c}_{i}$ is given by (24):
$d_{M}^{2}\left(\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,p});\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,0}),\boldsymbol{\tilde{\Sigma}}_{i}\right)=\boldsymbol{\Delta}_{i,p}^{T}\left(\boldsymbol{\tilde{\Sigma}}_{i}+\epsilon I^{(N-1)}\right)^{-1}\boldsymbol{\Delta}_{i,p},$ (113)
where as in (79), inversion has been empirically obtained with $\epsilon=10^{-6}$ and the N1N-1-dimensional identity matrix $I^{(N-1)}$. We again turn to the Schur complement Horn & Johnson (2013) and decompose (113):
$d_{M}^{2}\left(\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,p});\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,0}),\boldsymbol{\tilde{\Sigma}}_{i}\right)=\left(\boldsymbol{\Delta}_{i,p}^{(d)}\right)^{T}\left(\boldsymbol{\tilde{\Sigma}}_{i}^{(d)}+\epsilon I^{(d)}\right)^{-1}\boldsymbol{\Delta}_{i,p}^{(d)}+\boldsymbol{r}_{i,p}^{T}\boldsymbol{S}_{i}^{-1}\boldsymbol{r}_{i,p},$ (114)
with $\boldsymbol{r}_{i,p}\in\mathbb{R}^{m}$ and the Schur complement $\boldsymbol{S}_{i}\in\mathbb{R}^{m\times m}$ being:
$\boldsymbol{r}_{i,p}$ $=\boldsymbol{\Delta}^{(\perp)}_{i,p}-\boldsymbol{\tilde{C}}_{i}^{T}\left(\boldsymbol{\tilde{\Sigma}}_{i}^{(d)}+\epsilon I^{(d)}\right)^{-1}\boldsymbol{\Delta}^{(d)}_{i,p}$, (115)
$\boldsymbol{S}_{i}$ $=\boldsymbol{\tilde{\Sigma}}_{i}^{(\perp)}-\boldsymbol{\tilde{C}}_{i}^{T}\left(\boldsymbol{\tilde{\Sigma}}_{i}^{(d)}+\epsilon I^{(d)}\right)^{-1}\boldsymbol{\tilde{C}}_{i}.$ (116)
Let us define the set of squared Mahalanobis distances of cluster $\tilde{c}_{i}$ in dimension N1N-1 as:
$\mathcal{G}_{i}=\left\{d^{2}_{M}\left(\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,p});\boldsymbol{\Psi}_{t}(\mathbf{x}_{i}),\widetilde{\boldsymbol{\Sigma}}_{i}\right)\;\middle|\;p=1,\dots,N_{p}\right\},$ (117)
$d_{M}^{2}\left(\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,p});\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,0}),\boldsymbol{\tilde{\Sigma}}_{i}\right)-d_{M}^{2}\left(\boldsymbol{\Psi}^{(d)}_{t}(\mathbf{x}_{i,p});\boldsymbol{\Psi}^{(d)}_{t}(\mathbf{x}_{i,0}),\boldsymbol{\tilde{\Sigma}}^{(d)}_{i}\right)$= (118)
$\boldsymbol{r}_{i,p}^{T}\boldsymbol{S}_{i}^{-1}\boldsymbol{r}_{i,p}:=\delta_{\mathcal{G}_{i},p},$
$\mu_{\mathcal{G}_{i}}-\mu_{\mathcal{G}^{(d)}_{i}}=\frac{1}{|\mathcal{G}_{i}|}\sum_{g\in\mathcal{G}_{i}}g-\frac{1}{\Big{|}\mathcal{G}^{(d)}_{i}\Big{|}}\sum_{g\in\mathcal{G}^{(d)}_{i}}g=\frac{1}{N_{p}}\sum_{p=1}^{N_{p}}\boldsymbol{r}_{i,p}^{T}\boldsymbol{S}_{i}^{-1}\boldsymbol{r}_{i,p}=$ (119)
Similarly, we can express the deviation of the variance:
and with (119) and the Cauchy-Schwartz inequality, we can obtain:
and their deviations can be bounded by considering (119), (121):
with universal constants $C_{1},C_{2}>0$. Recalling (35), let the squared Mahalanobis distance from the output embedding to the cluster be:
$d^{2}_{M}\left(\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i}),\boldsymbol{\tilde{\Sigma}}_{i}^{(d)}\right):=a_{i},$ (126)
and employing (118) for the output embedding yields:
$d_{M}^{2}\left(\boldsymbol{\Psi}_{t}(\hat{\mathbf{x}}_{i});\boldsymbol{\Psi}_{t}(\mathbf{x}_{i,0}),\boldsymbol{\tilde{\Sigma}}_{i}\right)-d_{M}^{2}\left(\boldsymbol{\Psi}^{(d)}_{t}(\hat{\mathbf{x}}_{i});\boldsymbol{\Psi}^{(d)}_{t}(\mathbf{x}_{i,0}),\boldsymbol{\tilde{\Sigma}}^{(d)}_{i}\right)=\boldsymbol{r}_{i,a}^{T}\boldsymbol{S}_{i}^{-1}\boldsymbol{r}_{i,a}:=\delta_{\mathcal{G}_{i},a}.$ (127)
$\text{PM}_{i}=Q\left(k_{i},\frac{a_{i}}{\theta_{i}}\right).$ (128)
Consider the truncation-induced ellipsoid:
$\Bigl{|}\mathrm{PM}_{i}-\mathrm{PM}_{i}^{(d)}\Bigr{|}\;\leq\;\max_{(k_{c},\theta_{c},a_{c})\in\partial\mathcal{B}_{i}}\Big{|}\,Q\left(k_{c},a_{c}/\theta_{c}\right)-Q\left(\hat{k}^{(d)}_{i},\hat{a}^{(d)}_{i}/\hat{\theta}^{(d)}_{i}\right)\Big{|}$. (133)
$R_{i}=\max_{g\in\mathcal{G}_{i}}\,g.$ (134)
Let us consider the local box of values:
$\bigg{|}\mathrm{PM}^{(d)}_{i}-\widehat{\mathrm{PM}}^{(d)}_{i}\bigg{|}\leq\max_{(k_{c},\theta_{c},a_{c})\in\partial\mathcal{B}_{i}^{\mathrm{loc}}}\Big{|}\,Q\left(k_{c},a_{c}/\theta_{c}\right)-Q\left(\hat{k}^{(d)}_{i},\hat{a}^{(d)}_{i}/\hat{\theta}^{(d)}_{i}\right)\Big{|}$. (144)
[…]
In the deterministic term, large cross-block coupling $\widetilde{\boldsymbol{C}}_{i}$ or residual spread $\widetilde{\boldsymbol{\Sigma}}^{(\perp)}_{i}$ again directly inflate the error radius via the Schur complement.
Appendix C Error Radius and Probabilistic Confidence Bound of the PCC and SRCC
In this Appendix, we propagate the frame-level error radius and probabilistic widths developed in Appendix B.1 and  B.2 to the reported PCC and SRCC values. e.g., as reported in Table 4.
where:
where:
and $\mathbb{E}_{\boldsymbol{\pi}}\left(\zeta^{q,l,\mathcal{P}}_{i,f}\right)=0$. Thus, the two-sided probabilistic half-width $p^{q,l,\mathcal{P}}_{i,f}\geq 0$ can be interpreted as:
$\mathbb{P}_{\boldsymbol{\pi}}\left(\Big{|}\varepsilon^{q,l,\mathcal{P}}_{i,f}-\mu^{q,l,\mathcal{P}}_{i,f}\Big{|}\leq p^{q,l,\mathcal{P}}_{i,f}\right)\geq 1-\delta^{\mathcal{P}}:=c^{\mathcal{P}},$ (151)
$\sigma^{q,l,\mathcal{P}}_{i,f}=\frac{c}{z_{c^{\ast}}},$ (152)
We now propagate these errors from frame to utterance level, based on the aggregations we introduced in (37) and (41). On average, experiments showed that frames more than $g=4$ apart are effectively independent both for speech and music mixtures. Given the set $\mathcal{F}^{l}$ of time frames with two or more active sources, the standard Bartlett block-decimation Bartlett (1946) yields the conservative inflation:
$s(u)=0.999+4\bigl{(}1+\exp(-1.3669\,u+3.8224)\bigr{)}^{-1}.$ (156)
Let $W$ be the window and $H$ the hop of frame used for aggregation, then $M^{l}$ is the maximal number of possible windows. By norm submultiplicativity and the mean-value theorem (Horn & Johnson, 2013, Sec. 5.6):
and similarly, denoting $\hat{v}^{q,l,\mathcal{P}}_{i,\mathcal{A}}$ as the estimated aggregated measure across systems, where $\mathcal{A}$ is either average or PESQ-like aggregation (4.2), then its integration is given by:
For every vector $\mathbf{v}$, we denote its centered version by $\tilde{\mathbf{v}}$. Let us denote the PCC value between an observation vector $\mathbf{v}$ and a MOS vector $\mathbf{m}$ as $r^{\mathrm{PCC}}\left(\mathbf{v},\mathbf{m}\right)$, according to (45) and (46). Its gradient with respect to $\mathbf{v}$ at point $\hat{\mathbf{v}}^{l,\mathcal{P}}_{i,\mathcal{A}}$ is Benesty et al. (2009):
Then, the induced PCC bias can be bounded by:
$\boldsymbol{\eta}\sim\mathcal{N}\!\left(\mathbf{0},\mathrm{diag}\left(\left(\frac{h^{1,l,\mathcal{P}}_{i,\mathcal{A}}}{z_{c^{\ast}}}\right)^{2},\ldots,\left(\frac{h^{Q,l,\mathcal{P}}_{i,\mathcal{A}}}{z_{c^{\ast}}}\right)^{2}\right)\right),$ (164)
Turning to the SRCC, let $\rho^{\mathrm{SRCC}}(\cdot,\cdot)$ denote Spearman’s rank correlation between two vectors Kendall & Gibbons (1990), as defined in (47) and (48). Because ranks are piecewise-constant, a safe deterministic error radius is obtained by checking the two extreme bias orientations:
$h^{l,\mathcal{P},\textrm{SRCC}}_{i,\mathcal{A}}=\;\mathrm{Quantile}_{c^{\ast}}\!\left(\left|\rho^{\mathrm{SRCC}}\big{(}\hat{\mathbf{v}}^{l,\mathcal{P}}_{i,\mathcal{A}}+\boldsymbol{\eta},\mathbf{v}^{l}_{i,\mathrm{MOS}}\big{)}-\rho^{\mathrm{SRCC}}\big{(}\hat{\mathbf{v}}^{l,\mathcal{P}}_{i,\mathcal{A}},\mathbf{v}^{l}_{i,\mathrm{MOS}}\big{)}\right|\right),$ (167)
where we used $10^{4}$ draws for estimation, in the spirit of quantile bootstrap Efron & Tibshirani (1994).
To yield the probabilistic term, we assume that within any fixed trial $l$, the pairwise correlation between the source jitters has been empirically estimated and is denoted by $\rho_{l}$, while jitters from different trials are independent. This assumption holds by the construction of our trials in every scenario 4.1. Consequently, the $c$-level probabilistic half-width on the scenario mean equals:
$h^{\textrm{PCC}}\;=\;$ (170)
$h^{\textrm{SRCC}}\;=\;$ (171)
Appendix D Perceptual Distortions and Their Parameters Applied to the PS and PM
Table 6: Distortion bank used for the PM and PS measures. $f_{\textrm{s}}$ is the sampling frequency, ${A}_{95}$ marks the 95th-percentile absolute amplitude, and $A_{\mathrm{RMS}}$ the RMS amplitude of their respective signals.
@@ Appendix D Perceptual Distortions and Their Parameters Applied to the PS and PM
Appendix E Temporal and Spectral Examples of the Measures
Appendix F Frame Counts used for Normalized Mutual Information Calculations

### v2

@@ Appendix A Perceptual Distortions Applied in the PS and PM Calculations
Appendix B Additional Expreimental Setup Details
B.1 The PS and PM Measures
[…]
The empirical distributions of the frame-level values of the measures are shown in Figure 3. The PM and PS metrics exhibit contrasting distribution patterns. PM values cluster predominantly around zero with minimal density near one, while PS concentrate near one with virtually no occurrence near zero. Although frame-level human speech quality ratings are not publicly available for direct comparison, these patterns raise comparisons to how humans might perceive audio disturbances. The PM distribution aligns intuitively with human perception, as listeners typically penalize speech quality severely when disturbances occur, making ratings near the scale minimum unsurprising. However, real granular human ratings would likely show less extreme clustering around zero due to perceptual and rating scale complexities. The PS behavior presents a more complex interpretative challenge. Previous research suggests that humans perceive leakage as more quality-degrading than self-distortions, particularly in acoustic echo cancellation contexts (Khanagha et al., 2024), yet our findings here do not support this hypothesis. Whether this discrepancy stems from dataset characteristics, limitations of the PS measure itself, or the mismatch between granular PS values and aggregated human ratings remains unclear and warrants future investigation beyond the scope of this study.
B.2 The SEBASS Database
The SEBASS dataset suits this study for several reasons. Multilingual coverage of English and Spanish validates language-agnostic behavior, while music tests robustness to highly transient material. Large algorithmic spread creates rich output clusters that stress-test our methodology, and the dense sampling of raters allows for a more reliable estimation of the true mean-opinion score of subjective human opinion. Figure 4 shows that the speech reference signals have been recorded in a relatively clean environment with SNRs between 3.9 dB to 41.7 dB, with an average of 25 dB.
B.3 Pre-processing
We recognize that English and Spanish speakers rarely participate in the same conversation in real-life scenarios. To emulate realistic scenarios, we separate each 4-speaker mixture into their English and Spanish speakers, creating for each language two mixtures where the one has a pair of male speakers and the other a pair of female speakers. We acknowledge the uncertainty this step induces, as residuals of English may be present in the output signal of a Spanish speaker, and vice versa. It should be mentioned that listening tests have rendered this cross-language leakage extremely negligible. This may be since, as expected, source separation systems are able to leverage languages as a meaningful feature to recognize leakage and remove it.
Every waveform, including references, distortions, and outputs from all sources of the mixture, undergoes independent loudness normalization. We use the EBU Recommendation R-128 (European Broadcasting Union (EBU), 2011) and set the target level of each waveform to loudness units relative to full scale (LUFS) of -23. If the peak magnitude of the scaled waveform exceeds one, we attenuate it to avoid digital clipping. This step removes loudness bias, known to wrongly affect both human and algorithmic quality judgments, while preserving inter-speaker level relations across the outputs. Since the PS and PM measures address source separation, we also filter out any frames in which there are not at least two active sources using energy-based thresholding.
When applying diffusion maps, we set $\alpha=1$ in (3) to eliminate density-dependent bias from the embedding. This choice ensures that the PS and PM measures reflect the intrinsic geometric structure of the manifold rather than sampling density variations, which would introduce instead artificial distortions into the representation. We set $t=1$ in (6), to keep the diffusion operator focused on local neighborhoods and not being blurred by multi-step mixing.
B.4 Frame-level to Utterance-level Aggregation of the Measures
At trial $l$, let us denote the PM value of the $i$-th output of source-separation system $q$ in time frame $f$ as $\textrm{PM}^{q,l}_{i,f}$. Let $\mathcal{F}^{l}$ holds the time-frame indices with at least two active sources, and let $\mathcal{F}^{l}_{i}\in\mathcal{F}^{l}$ be its subset of time-frame indices in which the $i$-th source is active. Then, the utterance-level PM measure after average aggregation is given by:
Although average aggregation assumes that human listeners perceive global audio quality by weighing local events equally, which is evidently not the case (Rix et al., 2001), we chose to carry it for the PM since its behavior already exhibits strong and frequent granular penalties where the score drops to around zero. Thus, it is assumed that standard human behavior that weighs negative experience heavily in the utterance-level score is implicitly carried out by the nature of the PM measure itself.
However, this is not the case for the PS measure. Here, the aggregation we applied is inspired by the window-based pooling and logistic mapping used inside PESQ (Rix et al., 2001). Again, considering only time frame indices in $\mathcal{F}^{l}_{i}$ and dropping the rest, let us consider a window of size $W$ frames that slides across the PS measure with a hop size of $H$ frames. Using the $p$-norm, we define the following:
We then calculate the following root mean square expression:
and eventually the aggregated PS measure is given by:
where the constants were chosen according to (ITU-T, 2007). Here, we penalize lower scores explicitly using the $p$-norm to better match human perceptual aggregation.
B.5 Correlation Coefficients Between Aggregated Measures and MOS
$\mathbf{v}^{l}_{i}=$ $\left(v^{1,l}_{i},\ldots,v^{Q,l}_{i}\right)^{T}.$ (52)
The PCC (Benesty et al., 2009) is measured twice, for the PS and the PM, as follows:
Let $\mathcal{R}:\mathbb{R}^{Q}\rightarrow R^{Q}$ be the ranking operator, which in the presence of ties assigns the average ranks. The SRCC (Sedgwick, 2014) is measured for the PS and the PM:
We report these correlation coefficients per English, Spanish, and music mixtures scenarios separately. Let us denote $N^{l}_{f}$ the number of active sources in trial $l$ during frame $f$. Then, given a scenario with $\mathcal{L}$ independent trials such that $l\in\{1,\ldots,\mathcal{L}\}$}, we mark the maximal number of sources in trail $l$ with $N^{l}_{\textrm{max}}$:
$N^{l}_{\textrm{max}}=\max_{f\in\mathcal{F}^{l}}{N^{l}_{f}}.$ (57)
Appendix C Additional Experimental Results
Architecture Checkpoint Scenario Transformer Layers
WavLM Large microsoft/wavlm-large English 24
WavLM Base microsoft/wavlm-base English 12
wav2vec 2.0 Large facebook/wav2vec2-large-lv60 English 24
wav2vec 2.0 Base facebook/wav2vec2-base English 12
HuBERT Large facebook/hubert-large-ll60k English 24
HuBERT Base facebook/hubert-base-ls960 English 12
wav2vec 2.0 Large facebook/wav2vec2-large-xlsr-53 Spanish 24
MERT m-a-p/MERT-v1-95M Music 12
We begin by analyzing how performance depends on the choice of the pre-trained self-supervised model, the purpose of which is encoding waveforms into perceptual representations before they are fed into the diffusion maps. Table 4 lists the models we examine in this study. We consider six different models for English mixtures, based on the wav2vec 2.0 (Baevski et al., 2020), WavLM (Chen et al., 2022), and HuBERT (Hsu et al., 2021) backbones, with Figure 5 demonstrating their layer-wise performance. When using “Large” versions of the models, for both PCC and SRCC values, earlier layers frequently produce representations that allow superior results that gradually decline toward deeper layers, showing approximately 10% average absolute degradation between extremes. Existing layer-wise analysis already reported that acoustic and phonetic content is richly represented in intermediate layers, while deeper layers shift toward semantic abstraction (Pasad et al., 2022; Vaidya and Kell, 2022). Additional work confirms that distortion sensitivity peaks in the lower or middle layers and diminishes in deeper ones (Tamm et al., 2023; Hung et al., 2022), and that pretrained models tend to lose low-level signal fidelity in their deepest layers (Moussa and Toneva, 2025). A notable data point appears in the final layers of wav2vec 2.0 with a sharp drop in performance, especially for PS. This is likely due to its contrastive learning pretraining objective, which drives later layers to specialize in predicting quantized latent codes rather than preserving acoustic detail. For the “Base” versions of the models, we observe a somewhat different behavior. At low and middle layers, their performance is often quite competitive with the “Large” variants, and in several cases the former even outperforms the latter in deeper layers. However, for WavLM, the gap widens toward the final layers, with the “Large” version consistently outperforming. Interestingly, wav2vec 2.0 Base does not exhibit the sharp degradation observed in its counterpart and instead its deeper layers remain stable and even show improvements for PS, suggesting that the absence of over-specialization to quantized prediction in the “Base” model preserves sensitivity to perceptual distortions.
Table 5 narrows these models down to their top performing layer, chosen by the max-min criteria of the PCC and SRCC values, across all layers. A no-encoding option is also reported, where waveforms are skip-connected directly into the diffusion maps, which under-performs compared to encoded modes and emphasizes the effectiveness of the waveform encoding in the proposed pipeline. These results reaffirm that shallow layers achieve optimal performance. Although point-by-point comparisons show that “Base” models perform comparably to, or occasionally slightly exceed “Large” models, applying the max-min criteria across the models in the table reveals that ‘Large” models are preferable when jointly optimizing for PS and PM. For once, wav2vec 2.0 Large achieves for the PM a PCC and SRCC differences from its “Base” counterpart of absolute 6% and 2%, respectively, even when the PS case shows a negligible gap. Among the “Large” model variants, wav2vec 2.0 Large with transformer layer 2 emerges as the ideal configuration and we carry it forward as a case study we investigate. It should be noted that among “Large” models, the PS very slightly changes with roughly 1% and 0.5% gaps between extremes for the PCC and SRCC, respectively, while the PM gaps are more meaningful. This suggests that the choice of model may mainly affect the PM scores.
Measure Representation Transformer Layer SRCC PCC
PS wav2vec 2.0 (Large) 2 84.12% 83.74%
PS wav2vec 2.0 (Base) 2 84.25% 83.23%
PM wav2vec 2.0 (Large) 2 84.69% 86.36%
PM wav2vec 2.0 (Base) 2 82.79% 80.07%
PS WavLM (Large) 3 84.80% 84.16%
PS WavLM (Base) 2 84.84% 84.19%
PM WavLM (Large) 3 85.71% 81.44%
PM WavLM (Base) 2 82.82% 77.51%
PS HuBERT (Large) 3 84.48% 83.09%
PS HuBERT (Base) 2 84.83% 82.73%
PM HuBERT (Large) 3 84.12% 82.24%
PM HuBERT (Base) 2 81.37% 79.47%
PS Waveform (raw) - 73.42% 71.04%
PM Waveform (raw) - 69.30% 66.62%
Next, we analyze all scenarios with wav2vec 2.0 “Large” encoders for speech and the MERT encoder for music representations, and analyze the effect of their transformer layer on performance. The results are shown in Figure 6. English and Spanish mixtures, both evaluated with wav2vec 2.0 backbones, show broadly similar trends across layers, with Spanish exhibiting a sharper decline in deeper layers. This can be explained by the XLSR pretraining data being relatively scarcer in Spanish than in English, leading later layers to emphasize cross-lingual abstractions over fine acoustic detail (Conneau et al., 2020). Music mixtures with drums show the lowest performance among scenarios, which we attribute to the dominance of strong percussive transients. Self-supervised models have demonstrated less stability in these highly non-stationary regions, reducing the ability of PS and PM to capture perceptual degradations (Zeghidour et al., 2021). In contrast, music mixtures without drums demonstrate consistently high performance, in most layers even surpassing speech mixtures. This likely stems from the MERT backbone being particularly suited in capturing harmonic and timbral structure, allowing the measures to remain faithful to perceptual cues such as instrument texture and vocal clarity (Li et al., 2023). Interestingly, whether drums are present or not, MERT-based performance demonstrates a steady behavior across all layers, suggesting the MERT representations are not vulnerable to degradation across processing stages. The max-min criteria across all layers, per scenario, shows that the ideal layers for English, Spanish, drums, and no-drums music mixtures are layers 2, 2, 1, and 3, respectively.
We employed these layers to construct Table 1 in the main text and now we extend the discussion on it. The advantage of PESQ can be attributed to its long-standing perceptual model, which explicitly encodes aspects of loudness perception, asymmetry, and time-alignment penalties, features that directly penalize separation artifacts. In Spanish mixtures, the PS and PM are most performant in terms of PCC, but fall behind PESQ and SDR-based metrics in SRCC. One possible explanation is that the syllable-timed rhythm and steady vowels of Spanish make fidelity-driven metrics such as SI-SDR, CI-SDR, and SDR more predictive of listener rankings, as these metrics emphasize reconstruction accuracy at the waveform level. For music mixtures, PS and PM achieve the strongest overall correlations across both drums and no-drums conditions for both PCC and SRCC. Even though SpeechBERTscore has shown impressive results and is also based on a self-supervised backbone, it is mostly not competitive with our measures, and notably even performing worse than our raw waveform version at times, which projects on the importance of the diffusion maps in the pipeline. We emphasize that unlike English, we only inspected one backbone model for Spanish or music mixtures. In addition, the aggregation strategies we applied were not data-driven but a heuristic and reasoning-based choice. Consequently, while the proposed measures already demonstrate strong alignment with human perception, these low-hanging fruits may potentially boost performance. Quite surprisingly, the first group of STOI, PESQ, and SDR-based measures is consistently preferable to the second group consisting of DNSMOS, speechBERTscore, UTMOS, and others, which rarely achieve more than 70% in performance. One crucial conclusion this table suggests is that measures originally developed for a certain audio application, should not be zero-shot adapted into other applications, and in that case into source separation evaluation. Otherwise, values that drift from human opinion may be reported, which may spiral the development of audio technologies instead of accelerating it.
An additional stress test for our measures concerns their robustness to temporal misalignment between the input and output streams of the separator, a phenomenon commonly introduced by modern communication systems or, e.g., when dealing with references obtained from different, per-speaker microphones, such as in meeting datasets (Carletta et al., 2005; Vinnikov et al., 2024). Figure 7 illustrates the effect of artificial delays ranging from 0 ms to 100 ms across English, Spanish, and music scenarios. While performance gradually degrades for speech scenarios as misalignment grows, as expected, a 20 ms delay or less still preserves coefficients higher than 80%. Surpassing this threshold, however, often causes a pronounced drop that underscores this weakness in our measures, since human ratings are insensitive to these short latencies. Music mixtures exhibit a different pattern, as performance remains largely stable across delays, with the presence of drums introducing more variability than its counterpart.
Table 6 provides complementary information to the NMI results in Figure 1 but listing the frame counts used for the PS and PM measures for every examined threshold. Even the lowest threshold of 0.1, had a minimum of 481 for calculations, rendering its results statistically reliable. An interesting observation in the music scenario, without drums, is that it exhibits significantly more time frames in which there are at least two active sources, compared to all other scenarios.